OpenStudy (milrawr):

(f • g)=82 ( circle is open) f(x)= log3 (x+5) g(x)=log3 (x-1)

OpenStudy (milrawr):

the 3 is a base

OpenStudy (sunnnystrong):

This is a composite function. f circle g means f composed with g. f(g(x)) f(x)= log3 (x+5) // g(x)=log3 (x-1) Than f(log3(x-1) --> log3(log3(x+5)-1) -----

OpenStudy (sunnnystrong):

log3(log3(x+5)-1) = 82 Rewrite solve for x --> 3^82 = log3(x+5)-1 3^82+1 = log3(x+5) 1.3 x 10^39 = log3(x+5) 3^(1.3x10^39) = x+5 x= 3^(1.3x10^39) - 5 x= approx. 1.3 x 10^(39)